OpenStudy (anonymous):

Find the last two digits of 6 raised to the power of 13.

OpenStudy (dan815):

alll wrong it is

OpenStudy (zzr0ck3r):

what class is this for?

OpenStudy (zzr0ck3r):

are you doing stuff with modulo 10?

OpenStudy (anonymous):

A prereq math class

OpenStudy (anonymous):

No module 10

OpenStudy (anonymous):

so it is not 36?

OpenStudy (dan815):

no!

OpenStudy (dan815):

6 is the last number

OpenStudy (dan815):

2nd last number trying to figure out for odd powers

OpenStudy (dan815):

96???

OpenStudy (anonymous):

Thank you

OpenStudy (dan815):

ya i think so too i did something messed up in there lol

OpenStudy (dan815):

lemme tell u what i did, tell me if i messsed up

OpenStudy (anonymous):

I am with a study group right now who also thinks it may be 16 now that it has been mentioned.

jimthompson5910 (jim_thompson5910):

6^0 = 1 (mod 100) 6^1 = 6 (mod 100) 6^2 = 36 (mod 100) 6^3 = 6^2*6^1 = 36*6 = 216 = 16 (mod 100) 6^4 = (6^3)*6^1 = 16*6 = 96 (mod 100) 6^5 = (6^4)*6^1 = 96*6 = 576 = 76 (mod 100) 6^6 = (6^5)*6^1 = 76*6 = 456 = 56 (mod 100) 6^7 = (6^6)*6^1 = 56*6 = 336 = 36 (mod 100) So to sum things up 6^0 = 1 (mod 100) 6^1 = 6 (mod 100) 6^2 = 36 (mod 100) 6^3 = 16 (mod 100) 6^4 = 96 (mod 100) 6^5 = 76 (mod 100) 6^6 = 56 (mod 100) 6^7 = 36 (mod 100) the pattern starts off with 1, 6, but then gets into a loop of 36, 16, 96, 76, 56 and this loop repeats forever

OpenStudy (johnweldon1993):

Aww comment got deleted somehow...aww lol...I'm trying to figure it out too!

OpenStudy (johnweldon1993):

Now that @jim_thompson5910 has shown that...it seems easy! lol

OpenStudy (dan815):

ohh its 16 i multiplied 36 in the end

OpenStudy (dan815):

i did it for 36^7 oops

OpenStudy (dan815):